if a,8,b are in AP and a,4,b are in GP a,x,b are in HP then x =
Answer by richwmiller(17219)

AM * HM = GM^2
8*4=32
GM=sqrt(32)
GM=sqrt*(16*2)
GM=4sqrt(2)